OpenStudy (anonymous):

What is the slope of the line whose equation is 2x – 4y = 10?

Mehek (mehek14):

first change it to y = mx + b

OpenStudy (abdullahm):

^

OpenStudy (anonymous):

how

OpenStudy (abdullahm):

Add 4y on both sides Subtract both sides by 10 Divide both sides by 4

Mehek (mehek14):

move 2x to the other side

OpenStudy (anonymous):

so what's the answer

Mehek (mehek14):

try to solve for y first

OpenStudy (anonymous):

idk how

OpenStudy (abdullahm):

Then you must figure it out :)

OpenStudy (abdullahm):

We cannot do the work for you. The rules simply do not allow it.

OpenStudy (anonymous):

the "rules" can smd

OpenStudy (anonymous):

I needs helps

OpenStudy (anonymous):

hhhhhhhheeeeeeeeeeeeelllllllllllllllllllllpppppppppppppp

OpenStudy (abdullahm):

Sorry, but you do not want to cooperate.

OpenStudy (abdullahm):

2x – 4y = 10 Add 4y to both sides 2x - 4y + 4y = 10 + 4y Can you simplify it a little more? :)

OpenStudy (anonymous):

idk how

OpenStudy (abdullahm):

Look at the left side -4y + 4y = ? What does it equal?

OpenStudy (anonymous):

0?

OpenStudy (abdullahm):

Correct!

OpenStudy (abdullahm):

So now we have 2x = 10 + 4y Subtract 10 on both sides 2x - 10 = 10 - 10 + 4y What is 10 - 10 = ?

OpenStudy (anonymous):

0

OpenStudy (abdullahm):

Correct, once again! See! You do know something! :)

OpenStudy (abdullahm):

So, now we have 4y = 2x - 10 We need to divide 4 on both sides What is 4/4 = ? What is 2/4 = ? And, finally, what is 10/4 = ?

OpenStudy (anonymous):

2.5?

OpenStudy (abdullahm):

Yes, 10/4 = 2.5 What about 4/4 = ? And 2/4 = ?

OpenStudy (anonymous):

1 & 0.5?

OpenStudy (abdullahm):

correct!

OpenStudy (abdullahm):

So now we have y = 0.5x + 2.5

OpenStudy (anonymous):

so what's the answer

OpenStudy (abdullahm):

Ok, now we got your equation in y = mx + b form :)

OpenStudy (anonymous):

now what?

OpenStudy (abdullahm):

In the form y = mx + b, the m is the slope :)

OpenStudy (abdullahm):

so in y = 0.5x + 2.5 what is the slope? Hint: it's the number before x

OpenStudy (anonymous):

0.5?

OpenStudy (abdullahm):

Correct! That's your answer! :D

OpenStudy (abdullahm):

Now, that wasn't so hard was it? :)

OpenStudy (anonymous):

but the answer is supposed to be a fraction?

OpenStudy (abdullahm):

oh, in that case it would be 1/2

OpenStudy (anonymous):

oh ok thank you so much for all your help

OpenStudy (abdullahm):

It was my pleasure! :D